OpenStudy (jkbo):

Please Help which of the following is the minimum value of the equation y=2x^2+5 A.0 B.2 C.5 D.-5

OpenStudy (zzr0ck3r):

when dealing with a quadratic, we know the min (when the x^2 terms Is positive) is given by x = -b/(2a) where a,b c are ax^2+bx+c

OpenStudy (zzr0ck3r):

or take the derivative. what class is this for?

OpenStudy (jkbo):

this is for alg.2

OpenStudy (zzr0ck3r):

ok use my first reply

OpenStudy (zzr0ck3r):

what does a=? what does b = ?

OpenStudy (jkbo):

a is 2 Im not sure about B.

OpenStudy (zzr0ck3r):

b is gone, so b = 0

OpenStudy (zzr0ck3r):

so min x = -b/(2a) = ?

OpenStudy (jkbo):

so its zero because 2(2) is 4 / 0 is 0

OpenStudy (zzr0ck3r):

correct, x =0 but we want the min VALUE, value=y we know the min is AT x=0 so plug in x=0 to the original function and what do you get?

OpenStudy (jkbo):

5?

OpenStudy (zzr0ck3r):

correct

OpenStudy (jkbo):

thank you

OpenStudy (zzr0ck3r):

so if we ask WHERE is the min value, we want an x if we ask WHAT Is the min value we want a y

OpenStudy (jkbo):

Thank you can you also help me solve one more problem?

OpenStudy (zzr0ck3r):

sure

OpenStudy (jkbo):

Identify the axis of symmetry of the parabola. A.x=5 B.x=4 C.x=6 D.X=3

OpenStudy (zzr0ck3r):

wait, first off. up there ^^^^ on the last question you wrote x=0 because 2*2 = 4/0 =0

OpenStudy (zzr0ck3r):

this is not why x=0 x = -b/(2a) = -0/(2*2) = -0/4 = 0/4 = 0

OpenStudy (zzr0ck3r):

4/0 does not make sense

OpenStudy (zzr0ck3r):

a = 2, b = 0 \[x=\frac{-b}{2a}= \frac{-0}{2*2}=\frac{0}{4}=0\]

OpenStudy (zzr0ck3r):

understand?

OpenStudy (jkbo):

yes

OpenStudy (zzr0ck3r):

ok onto the next equation

OpenStudy (zzr0ck3r):

question*

OpenStudy (zzr0ck3r):

its the line down the middle what x value is that?

OpenStudy (zzr0ck3r):

down the middle of the porabolla

OpenStudy (zzr0ck3r):

parabola*

OpenStudy (jkbo):

4

OpenStudy (zzr0ck3r):

yep

OpenStudy (jkbo):

would that be the answer?

OpenStudy (zzr0ck3r):

yeah

OpenStudy (jkbo):

Thank you